# Basic representations of genus zero nonabelian Hodge spaces

### v2

@@ 4.8. Painlevé VI
Let us make a few comments about all these Painlevé examples. Notice that, in several cases, the generic tree of the modified irregular class $\bm{\breve{\Theta}}$ breaks part of the symmetry of the equation, for example for Painlevé VI it breaks the symmetry of the 4 regular singular points, and for Painlevé III it breaks the symmetry between the two singularities. This is because the construction depends on the location of the point at infinity, and because to have a minimal number of Stokes circles in $\bm{\breve{\Theta}}$, we made the standard choice of having one of the singularities at infinity. On the other hand, in several cases, the generic tree makes apparent some symmetry that wasn’t directly visible from the fission forest: for example for Painlevé IV the generic tree makes it clear that there is a symmetry between the two Stokes circles at infinity and the two singularities at finite distance, and for degenerate Painlevé III, it makes clear that the forests of two nongeneric classes of representations coincide. The basic reason underlying these observations is that the different combinatorial invariants behave in a nice way with respect to different natural operations on connections: the fission forest for a nonmodified irregular class is invariant under Möbius transformations and twisting by any rank one connection, but not invariant under $\operatorname{SL}_{2}(\mathbb{C})$, while the diagram is invariant under $\operatorname{SL}_{2}(\mathbb{C})$, but not under Möbius transformations changing the location of infinity or twists other than those belonging to the $\operatorname{SL}_{2}(\mathbb{C})$ action. The takeaway from this is that, for any given nonabelian Hodge space, it can be useful to combine all these invariants to understand as many symmetries of the situation as possible.

### v3

@@ 4.9.4 Diagram
Let us make a few comments about all these Painlevé examples. Notice that, in several cases, the short fission tree breaks part of the symmetry of the equation, for example, for Painlevé VI it breaks the symmetry of the 4 regular singular points, and for Painlevé III it breaks the symmetry between the two singularities. This is because the construction depends on the location of the point at infinity, and because to have a generic representation with minimal rank, we need to have one of the singularities at infinity (which coincides with the standard choice). On the other hand, in several cases, the short fission tree makes apparent some symmetry that was not directly visible from the fission forest: for example, for Painlevé IV the generic tree makes it clear that there is a symmetry between the two Stokes circles at infinity and the two singularities at finite distance, and for degenerate Painlevé III, it makes clear that the forests of two nongeneric classes of nearby representations coincide. The basic reason underlying these observations is that the different combinatorial invariants behave in a nice way with respect to different natural operations on connections: the fission forest is invariant under Möbius transformations and twisting by any rank one connection, but not invariant under $\operatorname{SL}_{2}(\mathbb{C})$, while the diagram is invariant under $\operatorname{SL}_{2}(\mathbb{C})$, but not under Möbius transformations changing the location of infinity or twists with rank one connections having singularities at finite distance. The takeaway from this is that, for any given nonabelian Hodge space, it can be useful to combine all these invariants to understand as many symmetries of the situation as possible.
Acknowledgements
I am funded by the PNRR Grant CF 44/14.11.2022, “Cohomological Hall Algebras of Smooth Surfaces and Applications”. I also acknowledge the funding of FCiências.ID, the great working environment of the Grupo de Física Matemática, where most of this work was carried out, and Philip Boalch, Gabriele Rembado, Giordano Cotti, Giulio Ruzza, Davide Masoero, Gabriele Degano, Julian Barragán Amado, for useful discussions. I am also grateful to the referees for their useful comments which helped improve the quality of the manuscript.